# DEPF: A UAV Multispectral Object Detector with Dual-Domain Enhancement and Priority-Guided Mamba Fusion

### v1

@@ Introduction
The loss function of DEPF follows conventional object detection loss and classification loss. Therefore, the cross-entropy loss function (C. E. Shannon, 1948) is adopted for the optimizing of object detection task, while the Smooth L1 loss function (Ross Girshick, 2015) is adopted for the optimizing of regression task. Therefore, the loss function of DEPF could be described as:
@@
RetinaNet RGB 78.5 34.4 24.1 69.8 28.8 25.0 23.4
$\text{R}^{3}\text{Det}$ RGB 80.3 56.1 42.7 80.2 44.4 27.4 27.4
KFIoU RGB 76.3 50.0 33.2 78.2 35.7 54.7 28.0
RoITransformer RGB 61.6 55.1 42.3 85.5 44.8 61.6 32.9
YOLOV5 RGB 76.2 48.9 35.5 68.9 40.4 54.0 30.2
Oriented R-CNN RGB 80.1 53.8 41.6 85.4 43.3 60.8 32.7
PKINet-S RGB 76.4 53.0 40.8 79.1 45.7 59.0 31.2
@@
RetinaNet Infrared 88.8 35.4 39.5 76.5 32.1 54.5 30.4
$\text{R}^{3}\text{Det}$ Infrared 89.5 48.3 16.6 87.1 39.9 62.3 36.7
$\text{S}^{2}\text{ANet}$ Infrared 89.9 54.5 55.8 88.9 48.4 67.5 40.4
KFIoU Infrared 90.2 65.0 48.5 89.0 43.0 67.1 41.2
ReDet Infrared 89.8 53.9 43.3 84.8 33.2 61.0 36.2
GWD Infrared 89.8 39.6 26.5 75.0 23.2 50.8 30.7
PKINet-S Infrared 90.2 67.3 56.3 88.8 51.2 70.8 43.8
DTNet Infrared 90.2 65.7 78.1 89.2 67.9 78.2 52.9
@@
UA-CMDet RGB+Infrared 87.5 60.7 46.8 87.1 38.0 64.0 40.1
Halfway-Fusion RGB+Infrared 90.1 62.3 58.5 89.1 49.8 70.0 42.9
GLFNet RGB+Infrared 90.3 72.7 53.6 88.0 52.6 71.4 42.9
AR-CNN RGB+Infrared 90.1 64.8 62.1 89.4 51.5 71.6
MBNet RGB+Infrared 90.1 64.4 62.4 88.8 53.6 71.9
TSFADet RGB+Infrared 89.9 67.9 63.7 89.8 54.0 73.1
CIAN RGB+Infrared 90.1 63.8 60.7 89.1 50.3 70.8
$\text{C}^{2}\text{Former}$ RGB+Infrared 90.2 68.3 64.4 89.8 58.5 74.2 47.5
TarDAL RGB+Infrared 89.5 68.3 56.1 89.4 59.3 72.6 43.3
DMM RGB+Infrared 90.4 78.9 66.2 89.5 68.3 78.6 50.6
@@
Table 1: Comparative experiment results on the DroneVehicle dataset. ’-’ represents that the result is not reported.
Experiment
Datasets and Metrics
DEPF is evaluated on DroneVehicle (Sun et al., 2022) and VEDAI (Sebastien Razakarivony and Frederic Jurie, 2016) benchmarks.
DroneVehicle Dataset
@@
DEPF is built on MMdetection and MMrotate platform. All of the experiments are conducted on four NVIDIA RTX 3090 GPUs with 24GB of memory. The environment of training and testing is Pytorch 2.1.2 and CUDA 11.8. DEPF is trained with 12 epochs and batch size of 4. The AdamW optimizer is adopted with an initial learning rate of 0.0001 and a weight decay of 0.05. We adopt Vision-Mamba as our backbone network and $\text{S}^{2}\text{ANet}$ as our detection head. The sizes of input images are 512 $\times$ 640 for DroneVehicle dataset and 512 $\times$ 512 for VEDAI dataset respectively. For the settings of CSWM block, Haar is chosen as wavelet basis, and the level of 2D-DWT is set to 2.
Methods Car Truck Boat Van Others mAP@0.5
RetinaNet 48.9 16.8 4.4 5.9 14.1 18.0
YOLO-S 95.5 65.5 74.2 70.7 47.9 70.8
YOLOV3 91.3 51.1 65.2 62.6 30.7 60.2
ROITransfomer 77.3 56.1 56.7 60.2 42.8 58.6
DAGN 81.3 32.9 41.5 78.5 53.9 57.6
Faster-RCNN 71.6 49.1 35.6 57.0 29.5 48.6
$\text{S}^{2}\text{ANet}$ 73.0 39.2 13.9 32.3 23.1 36.3
$\text{C}^{2}\text{Former}$ 76.7 52.0 43.3 48.0 41.9 52.4
CMAFF 81.7 58.8 66.0 68.5 51.5 65.3
DMM 84.2 65.7 72.3 72.5 56.2 70.2
Ours 85.3 66.7 74.3 73.0 56.4 71.1
Table 2: Comparative experiment results on the VEDAI dataset, the best results are marked with bold font.
Methods Input Size Para (MB) $\downarrow$ FLOPs (G) $\downarrow$ FPS (img/s) $\uparrow$
TSFADet 512 $\times$ 640 104.7 109.8 19
$\text{C}^{2}\text{Former}$ 118.5 115.6 11
DMM 88.1 108.8 17
Ours 80.3 108.3 16
Table 3: The complexity comparison between our DEPF and SOTA on NVIDIA RTX 3090 GPU.
@@
Figure 5: The visual comparison results on the DroneVehicle datasets. The advantages of our method are marked with blue circles. DEPF overcomes incorrect classification, missed detections, and false detections in the low-light scenarios.
@@
CSWM FDR Z-oder PGMF 0.5 0.5:0.95
@@
Table 4: Ablation study on the designed module on the DroneVehicle dataset.
Comparing with SOTA Methods
Comparison Results on the DroneVehicle dataset
Table 1 shows the comparison results on the DroneVehicle dataset. Our DEPF is compared with single-modality methods and multi-modality methods respectively. It could be observed that DEPF has significant performance in terms of mAP@0.5 and mAP@0.5:0.95.
Concretely, for single RGB modality methods, RoITransformer (Ding et al. 2019) achieves best performance with mAP@0.5 of 61.6% and mAP@0.5:0.95 of 32.9%, which is surpassed by our DEPF by 17.3 points and 18.3 points respectively. Moreover, DEPF achieves obvious performance for the category detection, such as the result on Car with mAP@0.5 of 90.5% and Bus with mAP@0.5 of 90.0%. For single infrared modality methods, all of the detectors perform better than adopting RGB modality only due to the harmful effects of low-light RGB images. DEPF achieves best performance with mAP@0.5 of 78.9%, while surpassed by DTNet (Zhang et al. 2024) with mAP@0.5:0.95 of 52.9%. However, DEPF outperforms DTNet most in category detection, such as 3.7 points gains on Truck. DEPF also surpasses some of multispectral detectors. Comparing with $\text{C}^{2}\text{Former}$ (Yuan and Wei 2024), it obtains 4.7 points improvement with mAP@0.5 and 3.7 points improvement with mAP@0.5:0.95. Comparing with TarDAL, it obtains 7.3 points improvement with mAP@0.5 and 8.1 points improvement with mAP@0.5:0.95.
Figure 5 shows the visualization results on the DroneVehicle dataset. DEPF is compared with $\text{C}^{2}\text{Former}$ and DMM in the low-light scenarios, as the regions marked by blue circles show. In the first scene, $\text{C}^{2}\text{Former}$ detects a bus-stop as car and bus. DMM doesn’t have such false detection but it classifies the freight-car as car. By contrast, our DEPF avoids these mistakes. In the second scene, it could be observed that DEPF makes correct classification of a car at night. In the third scenarios, $\text{C}^{2}\text{Former}$ and DMM make false detection that regards oval flower beds as cars at night, while DEPF avoids this mistake. These results shows that our method achieve robustness detection and classification in the low-light scenarios.
Comparison Results on the VEDAI dataset
Table 2 shows the comparison results on the VEDAI dataset. It could be observed that DEPF outperforms many of detectors with mAP@0.5 of 71.1%. However, for the Car and Van, DEPF is surpassed by YOLO-S (Betti and Tucci 2023) with 95.5% and DAGN (Zhang et al. 2019b) with 78.5%. These two detectors only adopts RGB images to achieve object detection, they are be impacted by the redundant information from multi-modality fusion mechanism. Nevertheless, DEPF also surpasses them on Truck, Boat and Other categories. Comparing with challenging DMM (Zhou et al. 2025b), DEPF obtains 0.9 points improvement with mAP@0.5 and obtains most 2 points improvement with mAP@0.5 on Boat in terms of specific categories.
Comparison of Computational Complexity
Table 3 shows the comparison results of computational complexity between DEPF and SOTA methods. Notably, this experiment is conducted on a NVIDIA RTX 3090 GPU, and the size of test input tensor is 512 $\times$ 640 with batch size of 1. It could be observed that our method has 80.3 MB parameter count and 108.3 GFLOPs (Giga Floating-Point Operations Per second). Especially, comparing $\text{C}^{2}\text{Former}$, it reduces about 32.2% parameter count and 6.3% GFLOPs. Comparing with DMM, it only reduces 7.8 MB parameter count and 0.5 GFLOPs. However, DEPF has no advantage in terms of FPS. The inference speed of TSFANet is the highest with 19 img/s, while DEPF’s is 16 img/s, which is also lower than DMM. The reason of this result is that the the inference speed is slowed down by 2D-DWT and FFT in DDE module, they don’t have parameter count but has computational costs.
@@
Ablation Study
Ablation Study on the Designed Modules
@@
The ablation study about the designed modules is conducted and the result is shown in Table 4. Comparing model 1 and model 2, the mAP@0.5 and mAP@0.5:0.95 improve about 0.3 points and 1.1 points when Z-order fusion is replaced with PGMF module, which demonstrates that the priority scanning could achieve more effective fusion than Z-order scanning. Comparing model 3, model 4 and model 5, it could be seen that introducing CSWM and FDR block only could not improve the metrics. In contrast, he mAP@0.5 and mAP@0.5:0.95 improve about 0.5 points and 1 points when both of them are introduced. This result demonstrates that enhancing the brightness and recovering the texture-details are important for detection. After adding PGMF module, the metrics improve to the optimum, which demonstrates that priority scanning fusion with enhanced images feature could achieve more correct detection.
Visualization of fusion feature maps

### v2

@@ III-C Loss Function
The loss function of DEPFusion follows conventional object detection loss and classification loss. Therefore, the Focal Loss [29] function is adopted for the optimizing of object detection task:
$\mathcal{L}_{\det}=-\alpha_{t}\left[p\left(1-t\right)+t\left(1-p\right)\right]^{\gamma_{t}}\log\left(p_{t}\right)$ (31)
where $\alpha_{t}$ denotes the weight factor of class $t$, $p_{t}$ denotes the predicted value which represents the probability of class $t$.
The Smooth L1 loss [30] function is adopted for the optimizing of regression task:
[…]
where $y_{gt}$ and $y_{pred}$ denote the groundtruth value and predicted value respectively.
Therefore, the loss function of DEPFusion could be described as:
@@
TABLE I: Comparative experiment results on the DroneVehicle dataset. ’-’ represents that the result is not reported.
@@
RetinaNet [31] RGB 78.5 34.4 24.1 69.8 28.8 25.0 23.4
$\text{R}^{3}\text{Det}$ [32] RGB 80.3 56.1 42.7 80.2 44.4 27.4 27.4
KFIoU [33] RGB 76.3 50.0 33.2 78.2 35.7 54.7 28.0
RoITransformer [34] RGB 61.6 55.1 42.3 85.5 44.8 61.6 32.9
YOLOV5 [35] RGB 76.2 48.9 35.5 68.9 40.4 54.0 30.2
Oriented R-CNN [36] RGB 80.1 53.8 41.6 85.4 43.3 60.8 32.7
PKINet-S [37] RGB 76.4 53.0 40.8 79.1 45.7 59.0 31.2
@@
RetinaNet [31] Infrared 88.8 35.4 39.5 76.5 32.1 54.5 30.4
$\text{R}^{3}\text{Det}$ [32] Infrared 89.5 48.3 16.6 87.1 39.9 62.3 36.7
$\text{S}^{2}\text{ANet}$ [38] Infrared 89.9 54.5 55.8 88.9 48.4 67.5 40.4
KFIoU [33] Infrared 90.2 65.0 48.5 89.0 43.0 67.1 41.2
ReDet [39] Infrared 89.8 53.9 43.3 84.8 33.2 61.0 36.2
GWD [40] Infrared 89.8 39.6 26.5 75.0 23.2 50.8 30.7
PKINet-S [37] Infrared 90.2 67.3 56.3 88.8 51.2 70.8 43.8
DTNet [41] Infrared 90.2 65.7 78.1 89.2 67.9 78.2 52.9
@@
UA-CMDet [11] RGB+Infrared 87.5 60.7 46.8 87.1 38.0 64.0 40.1
GLFNet [42] RGB+Infrared 90.3 72.7 53.6 88.0 52.6 71.4 42.9
AR-CNN [43] RGB+Infrared 90.1 64.8 62.1 89.4 51.5 71.6 -
MBNet [44] RGB+Infrared 90.1 64.4 62.4 88.8 53.6 71.9 -
TSFADet [16] RGB+Infrared 89.9 67.9 63.7 89.8 54.0 73.1 -
CIAN [45] RGB+Infrared 90.1 63.8 60.7 89.1 50.3 70.8 -
$\text{C}^{2}\text{Former}$ [13] RGB+Infrared 90.2 68.3 64.4 89.8 58.5 74.2 47.5
TarDAL [8] RGB+Infrared 89.5 68.3 56.1 89.4 59.3 72.6 43.3
TFDet [46] RGB+Infrared 90.1 63.4 48.0 87.7 48.9 67.6 -
DMM [14] RGB+Infrared 90.4 78.9 66.2 89.5 68.3 78.6 50.6
AFFNet [47] RGB+Infrared 90.2 66.0 50.9 88.1 48.8 68.8 -
@@
IV Experiment
IV-A Datasets and Metrics
DEPFusion is evaluated on DroneVehicle [11] and VEDAI [48] benchmarks.
IV-A1 DroneVehicle Dataset
@@ IV-B Implementation Details
DEPFusion is built on MMdetection [49] and MMrotate [50] platform. All of the experiments are conducted on four NVIDIA RTX 3090 GPUs with 24GB of memory. The environment of training and testing is Pytorch 2.1.2 and CUDA 11.8. DEPFusion is trained with 12 epochs and batch size of 4. The AdamW optimizer is adopted with an initial learning rate of 0.0001 and a weight decay of 0.05. We adopt Vision-Mamba as our backbone network and $\text{S}^{2}\text{ANet}$ as our detection head. For the settings of CSWM block, Haar is chosen as wavelet basis, and the level of 2D-DWT is set to 2.
The sizes of input images are 512 $\times$ 640 for DroneVehicle dataset and 512 $\times$ 512 for VEDAI dataset respectively. Following official setting [11, 48], the DroneVehicle dataset are divided into 17990, 1469 and 8980 pairs of RGB-IR image as train, validation and test set, while the VEDAI dataset are divided into 971, 100, 175 pairs of RGB-IR image as train validation and test set. All of the groundtruth is annotated with Orientated Bounding Box (OBB) format.
@@
Figure 5: The visual comparison results on the DroneVehicle datasets. The advantages of our method are marked with blue circles. DEPFusion overcomes incorrect classification, missed detections, and false detections in the low-light scenarios.
TABLE II: The complexity comparison between our DEPFusion and SOTA on NVIDIA RTX 3090 GPU.
Methods Input Size Para (MB) $\downarrow$ FLOPs (G) $\downarrow$ mAP@0.5 $\uparrow$
TSFADet 512 $\times$ 640 104.7 109.8 73.1
$\text{C}^{2}\text{Former}$ 118.5 115.6 74.2
DMM 88.1 108.8 78.6
Ours 80.3 108.3 78.9
Figure 6: The visual comparison results on the VEDAI datasets. The weakness of SOTA methods are marked with red circles.
TABLE III: Comparative experiment results on the VEDAI dataset, the best results are marked with bold font.
Methods Modality Car Truck Tractor Camping Car Van Pick-up Boat Plane Others mAP50 mAP0.5:0.95
RetinaNet [31] RGB 48.9 16.8 15.9 21.4 5.9 37.5 4.4 21.2 14.1 20.7 8.5
YOLO-S [51] RGB 95.5 65.5 63.8 - 70.7 - 74.2 75 47.9 70.4 -
YOLOV3 [52] RGB 80.4 38.6 57.7 66.5 76.9 67.8 26.2 88.7 51.3 61.6 -
ROITransfomer [34] RGB 77.3 56.1 64.7 73.6 60.2 71.5 56.7 85.7 42.8 65.4 27.6
ROITransfomer [34] Infrared 76.1 51.7 51.9 71.2 64.3 70.7 46.9 83.3 28.3 24.6 60.5
Faster-RCNN [53] Infrared 71.6 49.1 49.2 68.1 57.0 66.5 35.6 71.6 29.5 55.4 21.6
Oriented R-CNN [36] Infrared 77.0 55.0 47.5 73.6 63.2 72.2 49.4 79.6 30.5 27.1 60.9
$\text{S}^{2}\text{ANet}$ [38] Infrared 73.0 39.2 41.9 59.2 32.3 65.6 13.9 12 23.1 40.0 17.2
$\text{C}^{2}\text{Former}$ [13] RGB+Infrared 76.7 52.0 59.8 63.2 48.0 68.7 43.3 47 41.9 55.6 22.9
CMAFF [54] RGB+Infrared 81.7 58.8 68.7 78.4 68.5 76.3 66.0 72.7 51.5 69.2 30.5
DMM [14] RGB+Infrared 84.2 65.7 72.3 79.0 72.5 78.8 72.3 93.6 56.2 75.0 39.7
Ours RGB+Infrared 85.3 66.7 74.1 80.9 73.0 79.4 74.3 94.1 56.4 76.0 40.6
IV-C Comparing with SOTA Methods
IV-C1 Comparison Results on the DroneVehicle dataset
Table I shows the comparison results on the DroneVehicle dataset. Our DEPFusion is compared with single-modality methods and multi-modality methods respectively. It could be observed that DEPFusion has significant performance in terms of mAP@0.5 and mAP@0.5:0.95.
Concretely, for single RGB modality methods, RoITransformer [34] achieves best performance with mAP@0.5 of 61.6% and mAP@0.5:0.95 of 32.9%, which is surpassed by our DEPFusion by 17.3 points and 18.3 points respectively. Moreover, DEPFusion achieves obvious performance for the category detection, such as the result on Car with mAP@0.5 of 90.5% and Bus with mAP@0.5 of 90.0%. For single infrared modality methods, all of the detectors perform better than adopting RGB modality only due to the harmful effects of low-light RGB images. DEPFusion achieves best performance with mAP@0.5 of 78.9%, while surpassed by DTNet [41] with mAP@0.5:0.95 of 52.9%. However, DEPFusion outperforms DTNet most in category detection, such as 3.7 points gains on Truck. DEPFusion also surpasses some of multispectral detectors. Comparing with $\text{C}^{2}\text{Former}$ [13], it obtains 4.7 points improvement with mAP@0.5 and 3.7 points improvement with mAP@0.5:0.95. Comparing with TarDAL, it obtains 7.3 points improvement with mAP@0.5 and 8.1 points improvement with mAP@0.5:0.95.
Fig 5 shows the visualization results on the DroneVehicle dataset. DEPFusion is compared with $\text{C}^{2}\text{Former}$ and DMM in the low-light scenarios, as the regions marked by blue circles show. In the first scene, $\text{C}^{2}\text{Former}$ detects a bus-stop as car and bus. DMM doesn’t have such false detection but it classifies the freight-car as car. By contrast, our DEPFusion avoids these mistakes. In the second scene, it could be observed that DEPFusion makes correct classification of a car at night. In the third scenarios, $\text{C}^{2}\text{Former}$ and DMM make false detection that regards oval flower beds as cars at night, while DEPFusion avoids this mistake. The fourth and fifth scene show DEPFusion’s accurate classification under extreme low-light condition. These results shows that our method achieve robustness detection and classification in the low-light scenarios.
IV-C2 Comparison Results on the VEDAI dataset
Table II shows the comparison results on the VEDAI dataset. It could be observed that DEPFusion outperforms many of detectors with mAP@0.5 of 71.1%. However, for the Car and Van, DEPFusion is surpassed by YOLO-S [51] with 95.5% and DAGN [55] with 78.5%. These two detectors only adopts RGB images to achieve object detection, they are be impacted by the redundant information from multi-modality fusion mechanism. Nevertheless, DEPFusion also surpasses them on Truck, Boat and Other categories. Comparing with challenging DMM [14], DEPFusion obtains 0.9 points improvement with mAP@0.5 and obtains most 2 points improvement with mAP@0.5 on Boat in terms of specific categories.
DEPFusion is evaluated on VEDAI dataset [48] with SOTA methods, the visualization results are shown in Fig 6. It could be observed that DEPFusion could achieve correct classification and detection. In the first scene, DEPFusion reduces mis-classification of the cars on the road, comparing with $\text{C}^{2}\text{Former}$. The second scene shows the false detection and mis-classification of SOTA methods. Our DEPFusion avoids these mistakes successfully. The similar performances could be observed in the last scene.
IV-C3 Comparison of Computational Complexity
Table III shows the comparison results of computational complexity between DEPFusion and SOTA methods. Notably, this experiment is conducted on a NVIDIA RTX 3090 GPU, and the size of test input tensor is 512 $\times$ 640 with batch size of 1. It could be observed that our method has 80.3 MB parameter count and 108.3 GFLOPs (Giga Floating-Point Operations Per second). Especially, comparing $\text{C}^{2}\text{Former}$, it reduces about 32.2% parameter count and 6.3% GFLOPs. Comparing with DMM, it only reduces 7.8 MB parameter count and 0.5 GFLOPs. For the accuracy, DEPFusion outperforms SOTA methods with mAP@0.5 of 78.9%, gaining 5.8 points, 4.7 points and 0.3 points respectively when it is compared with TSFADet, $\text{C}^{2}\text{Former}$ and DMM.
Figure 7: The comparison between original RGB images and enhanced RGB images, which are generated by the DDE module.
TABLE IV: Ablation study on the designed module on the DroneVehicle dataset.
@@ IV-C3 Comparison of Computational Complexity
CSWM FDR Z-order PGMF 0.5 0.5:0.95
@@
IV-C4 The Enhancement Effect of DDE Module
Fig 7 shows the enhancement effect of DDE module. Comparing with original RGB images, our enhanced RGB images obtained from DDE module could show the details in the dark regions. For example, on the right side of the first scene, the clustered cars are observed due to their clear details. Similarly, the clustered cars are observed clearly on the left side of the third scene. Therefore, DDE module enhances the global brightness and recovers the texture-details of low-light RGB images, which improves the complementarity in the multi-modality fusion stage.
@@ IV-C4 The Enhancement Effect of DDE Module
TABLE V: Ablation study of the hyper-parameter in the CSWM block. LF and HF represents low-frequency component and high-frequency component respectively.
Models Wavelet Basis Enhanced Objects mAP
Haar Sym2 LF HF Together 0.5 0.5:0.95
model 1 77.8 47.2
model 2 75.6 46.6
model 3 78.6 50.0
model 4 77.8 50.1
model 5 76.8 49.7
model 6 78.9 51.2
TABLE VI: Ablation study of scanning way in the CSWM block. CS and CSS represents Concatenate Scanning and Cross-Scale Scanning respectively.
Models CS CSS mAP@0.5 mAP@0.5:0.95
model 1 76.2 47.9
model 2 78.9 51.2
TABLE VII: Ablation study result on wavelet transformation level of 2D-DWT and kernel size combination in the CSWM block.
Levels of 2D-DWT mAP@0.5 mAP@0.5:0.95
1 76.3 48.5
2 78.9 51.2
3 78.9 50.9
4 78.7 50.6
Kernel Size Combination mAP@0.5 mAP@0.5:0.95
[5,7,9] 78.7 50.0
[3,5,9] 78.5 51.0
[3,7,9] 78.7 50.8
[3,5,7] 78.9 51.2
IV-D Ablation Study
IV-D1 Ablation Study on the Designed Modules
@@
The ablation study about the designed modules is conducted and the result is shown in Table IV. Comparing model 1 and model 2, the mAP@0.5 and mAP@0.5:0.95 improve about 0.3 points and 1.1 points when Z-order fusion is replaced with PGMF module, which demonstrates that the priority scanning could achieve more effective fusion than Z-order scanning. Comparing model 3, model 4 and model 5, it could be seen that introducing CSWM and FDR block only could not improve the metrics. In contrast, he mAP@0.5 and mAP@0.5:0.95 improve about 0.5 points and 1 points when both of them are introduced. This result demonstrates that enhancing the brightness and recovering the texture-details are important for detection. After adding PGMF module, the metrics improve to the optimum, which demonstrates that priority scanning fusion with enhanced images feature could achieve more correct detection.
IV-D2 Ablation Study on CSWM Module
The ablation study of the settings in CSWM block is conducted, as shown in Table V and Table VI. For the wavelet basis, adopting Haar wavelet improve the metrics more than adopting Sym2 wavelet, because the latter brings more noise of the local details. For the enhanced objects, only the LF takes part in the enhancement could achieve good effect, no matter what types of wavelet basis. This is because the HF contains less important texture-details about the images, introducing HF into enhancement brings more redundant information. In addition, only enhancing HF could not enhance the global brightness of images. Therefore, the texture-details are enhanced by our designed FDR block additionally.
Table VI shows the effect of designed Cross-Scale Scanning mechanism (CSS). Notably, the Concatenate Scanning (CS) means that combining simple Z-order sequences as input. From result, it could be seen that introducing CSS improves mAP@0.5 with 2.7 points and mAP@0.5:0.95 with 3.3 points. This is because CSS learns multi-scale features and focuses on local brightness distribution between difference scales. Therefore, the global brightness is enhanced after the scanning and combination of Mamba.
The effect brought by the transformation level of 2D-DWT in CSWM is evaluated in Table VII. It could be observed that the values of two metric are approximate when level=2 and level=3, with mAP@0.5 of 78.9% and mAP@0.5:0.95 of 51.0%. The metric values decline as the level=4. Therefore, from the perspective of computational costs, the transformation level is set to 2 during training.
In addition, Table VII shows the effect of kernel size combination in the CSWM block. The multi-scale feature of low-frequency component is obtained by three different sizes of convolution kernel. The results show that the span of size variation brings small effects on mAP@0.5 and mAP@0.5:0.95. These small effects occur due to the low multi-scale complementary caused by obvious size differences.
IV-D3 Visualization of fusion feature maps